# Contextual Gesture: Co-Speech Gesture Video Generation through Context-aware Gesture Representation

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM '25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755140isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Computing methodologies Procedural animation
@@ 3.1. Contextualized Gesture Representation
(1) $\begin{split}\mathcal{L}_{\text{NCE}}&=-\frac{1}{2N}\sum_{i}\left(\log\frac{\exp{S_{ii}/\tau}}{\sum_{j}\exp{S_{ij}/\tau}}+\log\frac{\exp{S_{ii}/\tau}}{\sum_{j}\exp{S_{ji}/\tau}}\right),~\end{split}$
@@ 3.2. Speech-conditioned Gesture Generation
3.3. Structure-Aware Image Generation
@@
Instead of directly rendering the warped feature maps into RGB images, we first predict a low-resolution feature map of size 256×256×32256$\times$ 256$\times$ 32256 × 256 × 32. Multi-resolution edge heatmaps are generated and used as structural cues to refine the feature maps. After performing deformation and occlusion prediction at each scale using TPS (Zhao and Zhang, 2022), the edge heatmaps are fed into the generator. Specifically, we integrate these heatmaps into the fusion block using SPADE (Park et al., 2019) and the prediction of residuals are element-wise added to the warped feature maps, ensuring precise structural alignment.
@@ 4. Experiments
Dataset. We utilize PATS (Ginosar et al., 2019; Ahuja et al., 2020) for the experiments. It contains 84,000 clips from 25 speakers with a mean length of 10.7s, 251 hours in total. For a fair comparison, following the literature (Liu et al., 2022b; He et al., 2024) and replace the missing subject, with 4 speakers are selected (Noah, Kubinec, Oliver, and Seth). All video clips are cropped with square bounding boxes, centering speaks, resized to $\times 256$ 256. We defer the additional details in the Appendix.
@@
Baseline Methods. We benchmark Contextual Gesture against several co-speech gesture video generation methods: (1) ANGIE (Liu et al., 2022b), (2) S2G-Diffusion (He et al., 2024), (3) MM-Diffusion (Ruan et al., 2023), and (4) EchoMimicV2 (Meng et al., 2024). The first two are conventional optical-flow-based methods. MM-Diffusion is capable of achieving joint audio-visual generation. EchoMimicV2 is the most recent diffusion based speech-avatar animation model pretrained on large scale data.
Evaluation Results. We present quantitative evaluations in table 2. Our approach significantly outperforms existing methods in both gesture motion and video quality metrics. We provide qualitative evaluations in fig. 4. MM-Diffusion is not able to handle complex motion patterns, leading to almost static results. ANGIE and S2G-Diffusion struggle with local regions, such as the hands, due to its reliance on unsupervised keypoints for global transformations, which neglects local deformations. EchoMimicV2 lacks the background motion modeling and is only capable of presenting the aligned centered avatars in the middle. In addition, it fails to achieve diversified gestures. In contrast, our method demonstrates high-quality video generation, particularly in the facial and body areas. The alignment between gesture and speech is notably enhanced through our speech-content-aware gesture latent representation. For example, when the actor says ”90 joules,” he points to the screen, and he emphasizes phrases like ”so two ways” and ”in each case” by raising his hands.
@@
MM-Diffusion (Ruan et al., 2023) 67.56 4.32 0.65 0.11 - 77.65 4.14 14 days 600 sec
@@ C.2. Feature Representation
The video data within PATS dataset include personal identity information, and we strictly adhere to the data usage license ``CC BY - NC - ND 4.0 International,'' which permits non-commercial use.
@@ Appendix D Additional Implementation Details
For pixel-level motion generation, we utilize Thin Plate Splines (TPS) (Zhao and Zhang, 2022). Our framework tracks 116 keypoints (68 for the face and 48 for the body). The number of TPS transformations $K$ is set to 29, with each transformation utilizing $N=4$ paired keypoints. In accordance with TPS methodologies, both the dense motion network and occlusion-aware generators leverage 2D convolutions to produce $\times 64$ 64 weight maps for optical flow generation, along with four occlusion masks at various resolutions (32, 64, 128, and 256) to facilitate image frame synthesis.
@@ Appendix E Gesture Motion Generation
In this formulation, $\mathcal{L}_{mask}$ predicts the masked motion tokens $t_{i}$ at each time step $i$ based on the input audio and the special $[\mathrm{MASK}]$ token. Conversely, $\mathcal{L}_{res}$ focuses on learning from multiple quantization layers, where $t_{i}^{j}$ represents the motion token from quantizer layer $j$ and $t_{i}^{1:j-1}$ includes the tokens from preceding layers. We also feed the predicted tokens into the RVQ decoder for gesture reconstructions, with velocity and acceleration losses (Tevet et al., 2022; Siyao et al., 2022).
@@ Appendix H Time and Resource Consumption
In table 5, we present a comparison of training and inference times against existing baseline methods. For audio-gesture generation, our model's training time is comparable, albeit slightly slower, than that of ANGIE (Liu et al., 2022b) and S2G-Diffusion (He et al., 2024), primarily due to the inclusion of additional modules. However, it is considerably faster than MM-Diffusion (Ruan et al., 2023). Notably, our method excels in inference speed, outperforming all other baselines.
While the training of image-warping and image refinement requires a lot of time, our method leads to a substantial reduction in overall time and resource usage compared to MM-Diffusion and other stable-diffusion-based video generation approaches. Furthermore, the generative masking paradigm we employ significantly cuts down inference times when compared to diffusion-based models like S2G-Diffusion or the autoregressive generations in ANGIE.
@@
MM-Diffusion $\sim$14d Generation $\sim$9d + Super-Resolution $\sim$5d $\sim$600 sec
@@ Appendix I User Study Details
The TPS transformation accepts $N$ pairs of corresponding keypoints $(p^{\mathbf{D}}_{i},p^{\mathbf{S}}_{i})$ for $i=1,2,\ldots,N$ (referred to as control points) from a driving image $\mathbf{D}$ and a source image $\mathbf{S}$. It outputs a pixel coordinate mapping $\mathcal{T}_{tps}\left(\cdot\right)$, which represents the backward optical flow from $\mathbf{D}$ to $\mathbf{S}$. This transformation is founded on the principle that 2D warping can be effectively modeled through a thin plate deformation mechanism. The TPS transformation seeks to minimize the energy associated with bending this thin plate while ensuring that the deformation aligns accurately with the control points. The mathematical formulation is as follows:
[…]
@@
[…]

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ '25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755140isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer visionccs: Computing methodologies Procedural animation
@@ 3.1. Contextualized Gesture Representation
[…]
@@ 3.2. Speech-conditioned Gesture Generation
$3.$ Structure-Aware Image Generation
@@
Instead of directly rendering the warped feature maps into RGB images, we first predict a low-resolution feature map of size $256\times 256\times 32$. Multi-resolution edge heatmaps are generated and used as structural cues to refine the feature maps. After performing deformation and occlusion prediction at each scale using TPS (Zhao and Zhang, 2022), the edge heatmaps are fed into the generator. Specifically, we integrate these heatmaps into the fusion block using SPADE (Park et al., 2019) and the prediction of residuals are element-wise added to the warped feature maps, ensuring precise structural alignment.
@@ 4. Experiments
Dataset. We utilize PATS (Ginosar et al., 2019; Ahuja et al., 2020) for the experiments. It contains 84,000 clips from 25 speakers with a mean length of 10.7s, 251 hours in total. For a fair comparison, following the literature (Liu et al., 2022b; He et al., 2024) and replace the missing subject, with 4 speakers are selected (Noah, Kubinec, Oliver, and Seth). All video clips are cropped with square bounding boxes, centering speaks, resized to $256\times 256$. We defer the additional details in the Appendix.
@@
Baseline Methods. We benchmark Contextual Gesture against several co-speech gesture video generation methods: (1) ANGIE (Liu et al., 2022b), (2) S2G-Diffusion (He et al., 2024), (3) $M$-Diffusion (Ruan et al., 2023), and (4) EchoMimicV2 (Meng et al., 2024). The first two are conventional optical-flow-based methods. $M$-Diffusion is capable of achieving joint audio-visual generation. EchoMimicV2 is the most recent diffusion based speech-avatar animation model pretrained on large scale data.
Evaluation Results. We present quantitative evaluations in table 2. Our approach significantly outperforms existing methods in both gesture motion and video quality metrics. We provide qualitative evaluations in fig. 4. $M$-Diffusion is not able to handle complex motion patterns, leading to almost static results. ANGIE and S2G-Diffusion struggle with local regions, such as the hands, due to its reliance on unsupervised keypoints for global transformations, which neglects local deformations. EchoMimicV2 lacks the background motion modeling and is only capable of presenting the aligned centered avatars in the middle. In addition, it fails to achieve diversified gestures. In contrast, our method demonstrates high-quality video generation, particularly in the facial and body areas. The alignment between gesture and speech is notably enhanced through our speech-content-aware gesture latent representation. For example, when the actor says ”90 joules,” he points to the screen, and he emphasizes phrases like ”so two ways” and ”in each case” by raising his hands.
@@
$M$-Diffusion (Ruan et al., 2023) 67.56 4.32 0.65 0.11 - 77.65 4.14 14 days 600 sec
@@ C.2. Feature Representation
The video data within PATS dataset include personal identity information, and we strictly adhere to the data usage license ``$C$ BY - NC - ND 4.0 International,'' which permits non-commercial use.
@@ Appendix D Additional Implementation Details
For pixel-level motion generation, we utilize Thin Plate Splines (TPS) (Zhao and Zhang, 2022). Our framework tracks 116 keypoints (68 for the face and 48 for the body). The number of TPS transformations $K$ is set to 29, with each transformation utilizing $N=4$ paired keypoints. In accordance with TPS methodologies, both the dense motion network and occlusion-aware generators leverage 2D convolutions to produce $64\times 64$ weight maps for optical flow generation, along with four occlusion masks at various resolutions (32, 64, 128, and 256) to facilitate image frame synthesis.
@@ Appendix E Gesture Motion Generation
In this formulation, $\mathcal{L}_{mask}$ predicts the masked motion tokens $t_{i}$ at each time step $i$ based on the input audio and the special $\mathrm{MASK}]$ token. Conversely, $\mathcal{L}_{res}$ focuses on learning from multiple quantization layers, where $t_{i}^{j}$ represents the motion token from quantizer layer $j$ and $t_{i}^{1:j-1}$ includes the tokens from preceding layers. We also feed the predicted tokens into the RVQ decoder for gesture reconstructions, with velocity and acceleration losses (Tevet et al., 2022; Siyao et al., 2022).
@@ Appendix H Time and Resource Consumption
In table 5, we present a comparison of training and inference times against existing baseline methods. For audio-gesture generation, our model's training time is comparable, albeit slightly slower, than that of ANGIE (Liu et al., 2022b) and S2G-Diffusion (He et al., 2024), primarily due to the inclusion of additional modules. However, it is considerably faster than $M$-Diffusion (Ruan et al., 2023). Notably, our method excels in inference speed, outperforming all other baselines.
While the training of image-warping and image refinement requires a lot of time, our method leads to a substantial reduction in overall time and resource usage compared to $M$-Diffusion and other stable-diffusion-based video generation approaches. Furthermore, the generative masking paradigm we employ significantly cuts down inference times when compared to diffusion-based models like S2G-Diffusion or the autoregressive generations in ANGIE.
@@
$M$-Diffusion $\sim$14d Generation $\sim$9d + Super-Resolution $\sim$5d $\sim$600 sec
@@ Appendix I User Study Details
The TPS transformation accepts NN pairs of corresponding keypoints $(p^{\mathbf{D}}_{i},p^{\mathbf{S}}_{i})$ for $i=1,2,\ldots,N$ (referred to as control points) from a driving image $\mathbf{D}$ and a source image $\mathbf{S}$. It outputs a pixel coordinate mapping $\mathcal{T}_{tps}\left(\cdot\right)$, which represents the backward optical flow from $\mathbf{D}$ to $\mathbf{S}$. This transformation is founded on the principle that 2D warping can be effectively modeled through a thin plate deformation mechanism. The TPS transformation seeks to minimize the energy associated with bending this thin plate while ensuring that the deformation aligns accurately with the control points. The mathematical formulation is as follows:
(11) $\begin{split}\min&\iint_{\mathbb{R}^{2}}\left(\left(\frac{\partial^{2}\mathcal{T}_{tps}}{\partial x^{2}}\right)^{2}+2\left(\frac{\partial^{2}\mathcal{T}_{tps}}{\partial x\partial y}\right)^{2}+\left(\frac{\partial^{2}\mathcal{T}_{tps}}{\partial y^{2}}\right)^{2}\right)\,dxdy,\\ &\text{s.t.}\quad\mathcal{T}_{tps}(p_{i}^{\mathbf{D}})=p_{i}^{\mathbf{S}},\quad i=1,2,\ldots,N,\end{split}$
@@
(14) $\delta_{(0,0)}=\begin{cases}\infty,&\text{if }(x,y)=(0,0)\\ 0,&\text{otherwise}\end{cases},\quad\text{and}\iint_{\mathbb{R}^{2}}\delta_{(0,0)}(x,y)\,dxdy=1,$